# Faster Diffusion via Temporal Attention Decomposition

### v2

@@ Appendix H Evaluation on Text-Image Alignment
Table A8: Generation performance on DPG-Bench [20]. CA and SA represent cross-attention and self-attention, respectively. For PixArt, parameters are set to $m=15$ and $k=3$, whereas for SDXL, we utilize $m=10$ and $k=5$. Evaluation scores are obtained using mPLUG-Large [25] with predefined questions.
@@
The performance of Tgate in text-image alignment is assessed based on the protocol outlined by ELLA [20]. Table A8 shows the competitive performance of Tgate compared with the baseline across various evaluation dimensions, indicating its effectiveness.
Table A9: Frame consistency on the Open-Sora sample dataset [23]. The frame consistency is calculated based on the L2 distance between the adjacent frames. For OpenSora, $m=100$ and $k=3$, while in SVD, $m=10$ and $k=5$.
@@
         Tgate (CA)          31.162
         Tgate (CA, SA)          31.020
@@ Appendix I Evaluation on Frame Consistency
The performance of Tgate in video generation is evaluated based on the frame consistency. The L2 distance between different frames are obtained, where a smaller value indicates a smoother change between frames. As shown in Table A9, Tgate does not cause significant differences in this metric, confirming its effectiveness.
Appendix J Limitations
Visualization results show that the images created by Tgate are similar to those made without Tgate, with only minor differences. Although Tgate obtains a better FID score in some configurations, it might be difficult for users to spot the difference. Nevertheless, the main benefit of using Tgate is the improvement of efficiency during denoising. The potential improvement in image quality is a spin-off.
Appendix K Broader Impacts

### v3

@@ Appendix F Error Bar
Appendix G Additional Comparison
Table A8: Generation performance of ToMe (Bolya & Hoffman, 2023) and SSD-1B (Gupta et al., 2024) with and without with Tgate on MS-COCO-10k (Lin et al., 2014) dataset. The base mode for ToMe is SD-2.1(Rombach et al., 2022).
        FID         Latency
        ToMe         24.763         12.240s
        Tgate + ToMe         24.383         8.731s
        SSD-1B         28.564         30.097s
        Tgate         25.004         18.328s
We further incorporate our method into SSD-1B Gupta et al. (2024), a lightweight model distilled from a larger diffusion model. As shown in Table C, we test the models’ performance on the COCO Validation Set in terms of latency and FID-10k. Latency is defined as the time required to generate one image in a resolution of 768, and the results are collected on a platform using an Nvidia 1080ti. The results demonstrate that our method is also compatible with models distilled from larger models.
We also compare our method with ToMe Bolya & Hoffman (2023). We set ToMe’s merging ratio to 50% and use SD-2.1 as the base model. The model utilizes the DPM-Solver with 25 inference steps as the noise scheduler. Our method achieves a latency of 11.372 seconds per image, slightly outperforming ToMe. Given that our approach is orthogonal to ToMe, integrating TGATE with ToMe results in an improved latency of 8.731 seconds. This empirical study demonstrates the effectiveness of our method and its potential for broad application when combined with various acceleration techniques.
@@ Appendix H Evaluation on Text-Image Alignment
Table A9: Generation performance on DPG-Bench (Hu et al., 2024). CA and SA represent cross-attention and self-attention, respectively. For PixArt, parameters are set to $m=15$ and $k=3$, whereas for SDXL, we utilize $m=10$ and $k=5$. Evaluation scores are obtained using mPLUG-Large (Li et al., 2022) with predefined questions.
@@
Table A10: Generation performance of SD-2.1 (Rombach et al., 2022) with Tgate on T2I-Compbench (Huang et al., 2023). We set $m$ as 10 and $k$ as 5.
Color Shape Texture Spatial Non-Spatial Complex
SD-2.1 43.31 42.27 47.24 6.52 29.80 30.50
SD-2.1 + Tgate (CA) 41.97 41.16 45.28 5.73 29.55 29.45
SD-2.1 + Tgate (CA,SA) 40.72 40.01 44.22 5.14 29.32 29.44
The performance of Tgate in text-image alignment is assessed based on the protocol outlined by CLIP score Radford et al. (2021), ELLA (Hu et al., 2024) and T2I-Compbench (Huang et al., 2023). Table A9 and Table A10 shows the competitive performance of Tgate compared with the baseline across various evaluation dimensions, indicating its effectiveness. We provide visualization samples in Fig. A7. As noted in our limitations, increased inference acceleration can lead to greater deviations from the baseline. While these differences are often perceptually indistinguishable, in some cases, they may cause noticeable artifacts, such as the altered shape of the bowl or slight distortion of the bracelet in Fig. A7. Given the acceleration achieved, this trade-off is acceptable in most cases that do not demand exceptionally high-quality samples.
Figure A7: Generated samples from (a) PixArt, (b) Tgate with CA cache, and (c) Tgate with both CA and SA cache. The caption is derived from T2I-Coompbench.
Table A11: CLIP score of Tgate on MJHQ (Li et al., 2023) and MS-COCO (Lin et al., 2014) datasets on PixArt (Chen et al., 2024b) and SDXL (Podell et al., 2023).
     Method      CLIP Score
     SDXL      34.148
     Tgate ($m$=10)      33.371
     Tgate ($m$=10 $k$=5)      32.710
     PixArt-Alpha      34.233
     Tgate ($m$=15)      33.872
     Tgate ($m$=15 $k$=3)      33.673
Table A12: Frame consistency on Open-Sora sample dataset (Lab & etc., 2024). The frame consistency is calculated based on the L2 distance between the adjacent frames.
@@
Tgate ($m=10$) 31.162
Tgate ($m=10$,$k=5$) 31.020
Table A13: Memory overhead caused by Tgate. The base model is PixArt (Chen et al., 2024b) and SDXL (Podell et al., 2023) and the computational platform is a single V100 GPU card with pytorch 2.2.
      Memory Cost
      SDXL       8515 MB
      Tgate ($m$=10, $k$=5)       8531 MB
      PixArt-Alpha       13705 MB
      Tgate ($m$=15, $k$=3)       13707 MB
Figure A8: Generated samples based on SVD. SDXL generates the input frame with a caption of “A snowy forest landscape with a dirt road running through it. The road is flanked by trees covered in snow, and the ground is also covered in snow. The sun is shining, creating a bright and serene atmosphere. The road appears to be empty, and there are no people or animals visible in the video. The style of the video is a natural landscape shot, with a focus on the beauty of the snowy forest and the peacefulness of the road.”.
@@ Appendix I Evaluation on Frame Consistency
The performance of Tgate in video generation is evaluated based on the frame consistency. The L2 distance between different frames are obtained, where a smaller value indicates a smoother change between frames. As shown in Table A13, Tgate does not cause significant differences in this metric, confirming its effectiveness. We also provide a visualization in Fig. A8
Appendix J Evaluation on Memory Cost
We evaluate the memory overhead of Tgate, which accelerates inference by caching features and may increase GPU usage. As shown in Table A13, using SDXL and Pixart-Alpha as baselines for U-Net and transformer models, Tgate incurs only a minimal, single-digit memory cost, which is negligible in most cases, demonstrating its feasibility.
Appendix K Discussion on Anchor Feature.
This paper employs the average feature maps from both unconditional and conditional branches as the anchor feature for caching, following CFG. We provide visualization samples using only one branch instead of averaging. As shown in Fig. A9, the results suggest minimal impact on performance. This aligns with our pilot study (Fig. 2), where attention maps converge to a fixed point after several inference steps, including for the conditional/unconditional branches in CFG.
Figure A9: Generated samples of (a) PixArt and (b) Tgate reusing the averaged of cross-attention maps, (c) Tgate reusing the unconditional cross-attention maps, and (d) Tgate using text-conditional cross-attention maps.
Appendix L Hyper-parameter Selection
Tgate is a simple method for training-free acceleration with two hyperparameters, $k$ and $m$. Their values can adapt to different inference steps to achieve a balanced trade-off. As shown in Fig. A10, $m$ and $k$ are set as $3/5$ and $1/5$ of the total inference steps, respectively. Increasing inference steps minimally impacts the generated results, suggesting that $k$ and $m$ can be reliably set as fixed proportions of the inference steps for most scenarios.
[…]
Appendix M Details of Prompt
We showcase all prompts used in this paper to ensure reproducibility.
Table A14: Figures and Corresponding Prompts
Figure Caption
Figure 1 ’A desert oasis with a mirage of a city.’
’A galaxy with colorful nebulae and star clusters.’
’An underwater coral reef with mermaids and treasure.’
’A magical duel in a ruined castle.’
’A painting of mountains, in the style of Monet.’
Figure 3 ’High quality photo of an astronaut riding a horse in space.’
Figure 4 ’Paisaje montañoso nevado.’
Figure 6 (a) ’Beautiful lady, freckles, big smile, blue eyes, short ginger hair, dark makeup, wearing a floral blue vest top, soft light, dark grey background.’
’Professional portrait photo of an anthropomorphic cat wearing fancy gentleman hat and jacket walking in autumn forest.’
’Stars, water, brilliantly, gorgeous large scale scene, a little girl, in the style of dreamy realism, light gold and amber, blue and pink, brilliantly illuminated in the background.’
’A legendary figure in Cyberpunk City, rainy night, front, sunglasses.’
Figure 6 (b) ’Woman standing in front of a window with her hair blowing, modern anime, sunshine, highly detailed.’
’Create a story or description inspired by this photo of a gray-haired 24 year old Norwegian woman in front of the sun with subtle freckles and hazel eyes.’
’A layer cake made out of a stratographic cross-section of the Sonoran Desert.’
’A teal flower in a barren garden, beautiful villa background, octane, redshift, highly detailed.’
Figure A4 ’a delicious salad with beef, stock photo.’
Figure A5 ’Sunset over a tranquil lake in the mountains’
’Neon-lit skyline of a bustling future metropolis. ’
’A coral reef bustling with diverse marine life.’
’Gears and steam power an old-world cityscape.’
’Dusty trails leading to a desert sunset.’
’Red hot air balloon safari at sunrise’
’Giant flowers under a moon sky’
Figure A6 ’Self-portrait oil painting, a beautiful cyborg with golden hair, 8k.’
Figure A8 ’A snowy forest landscape with a dirt road running through it. The road is flanked by trees covered in snow, and the ground is also covered in snow. The sun is shining, creating a bright and serene atmosphere. The road appears to be empty, and there are no people or animals visible in the video. The style of the video is a natural landscape shot, with a focus on the beauty of the snowy forest and the peacefulness of the road.’
Figure A7 ’A cubic ice cube and a spherical ice bucket.’
’The blue bowl was on top of the white placemat.’
’The green plant was on the right of the white wall.’
’A diamond brooch and a teardrop bracelet’
Figure A10 ’a chicken on the right of a suitcase.’
"professional portrait photo of an anthropomorphic cat wearing fancy gentleman hat and jacket walking in autumn forest."
’stars, water, brilliantly, gorgeous large scale scene, a little girl, in the style of dreamy realism, light gold and amber, blue and pink, brilliantly illuminated in the background.’